OpenStudy (frankthabunnyy):

Find the domain of fg f(x) = x^2 +1 g(x) = 1/X

OpenStudy (solomonzelman):

\(\large\color{black}{ f(x)=x^2+1 }\) \(\large\color{black}{ g(x)=\frac{\Large{1}}{\Large{x}} }\) \(\large\color{black}{ f(g(x))=(\frac{\Large{1}}{\Large{x}})^2+1=\frac{\Large{1}}{\Large{x^2}} +1 }\)

OpenStudy (solomonzelman):

this is the g(x). You know that the denominator can't equal zero, so thus x can't equal zero. Any other number works for the domain.....